# Experimental Study of Underwater Acoustic Reconfigurable Intelligent Surfaces with In-Phase and Quadrature Modulation

### v1

@@ V-A2 Circuit Design
The load network comprises three subnetworks: a resistor network, a capacitive network, and an inductive network. The resistor network has only one stage designed to generate the in-phase or antiphase component required for IQ modulation. Conversely, both the capacitive and inductive networks feature three stages each, dedicated to producing the quadrature component for IQ modulation.
Resistor network contains an AD8403AR50 programmable potentiometer, whose resistance is managed by the MCU via the serial peripheral interface (SPI) protocol. Let $R_{0}$ represent the resistance of this potentiometer. According to the reflection coefficient presented in (1), when the load connected to the matching circuit is non-reactive and $Z_{0}$ is a real number, $\Gamma$ has no imaginary component. In this circumstance, the magnitude of the reflection coefficient varies between 0 and 1, while the phase shift of the reflected wave remains fixed at 0° for $R_{0}\geq Z_{0}$, or 180° for $R_{0}<Z_{0}$.
According to (1), when $Z_{0}$ is a real number, the load impedance must be a complex number to generate the quadrature component necessary for IQ modulation. In this scenario, it is impossible to continuously adjust the magnitude of the reflection coefficient by simply altering the resistance of the load impedance. To tackle this issue, in our design, the capacitive network includes three stages, designated as $Z_{1}$ to $Z_{3}$ in Fig. 5. Connecting different stages of this network to the matching circuit yields a reflection coefficient with a consistent phase of --90° and three selectable amplitudes increasing stepwise from 0.3 to 0.9. Conversely, $Z_{4}$ to $Z_{6}$ form a three-stage inductive network that produces a reflection coefficient with a consistent phase of 90°, and three optional amplitudes ranging from 0.3 to 0.9.
With the architecture depicted in Fig. 5, the MCU can effectively manipulate both the phase and magnitude of the reflection coefficient by adjusting the resistance of the potentiometer and the ON/OFF status of the NMOS transistors in each stage of the load network. In the following section, we will introduce the selection of $Z_{0}$ and discuss the stratege for managing the high dynamic range of the acoustic signal in the circuit design.
V-B High-Impedance Matching
@@
In order to precisely control the phase and amplitude of the quadrature component for accurate IQ modulation, the wiper resistance of programmable potentiometers cannot be ignored. For instance, models such as the AD8403 [26] and AD5245 [27] typically exhibit a wiper resistance of about 50 $\Omega$, while the MAX5427 has a resistance of 100 $\Omega$ [28]. Consequently, the minimum resistance of the load network in (1) does not approach zero but remains significantly higher. In scenarios where the impedance of the reflector is matched to 50 $\Omega$ (i.e., $Z_{0}\!=\!50$$\Omega$), the reflection coefficient consistently remains positive. As a result, the reflector will be incapable of generating an antiphase wave.
@@
V-C Highly Dynamic Incident Waves
@@
VI Performance Evaluation
This section evaluates the performance of the proposed UA-RIS through experiments. Initially, we examine the power consumption of the system. Subsequently, the results collected from both tank and lake environments are presented to demonstrate the viability of the UA-RIS in real-world applications.
VI-A Power Consumption
Although the UA-RIS does not transmit acoustic signals, it still requires an energy supply to maintain the operation of active components for IQ modulation. In our implementation, the power consumption of the system primarily originates from four components: the MCU, twelve I/O extenders, and six programmable potentiometers. In this section, we thoroughly analyze the power consumption of UA-RIS operating in various states.
@@ VI-A Power Consumption
VI-B Tank Experiments
VI-B1 Experimental settings
@@
VI-B2 Resistive load modulation
@@
VI-B3 Reactive load modulation
The multipath model presented in (6) can also be used to validate the reflector’s ability to produce the quadrature component for IQ modulation. This can be achieved by programming the MCU to sequentially connect the subnetwork shown in Fig. 5 with the acoustic reflector. The waveform of the received signal is depicted in Fig. 11 (a) [^1] The experiment in Fig. 11 was conducted on a different day than those in Figs. 8 to 10. This difference resulted in varying acoustic environments and channel states, indicating that this experiment should be considered independent., where the source is 28.23 kHz CW with 20 V peak-to-peak voltage. In the figure, labels such as C0.3 and L0.9 indicate that the acoustic reflector is connected to the capacitive and inductive load network with reflection coefficients of 0.3 and 0.9 amplitude, respectively.
@@ VI-B3 Reactive load modulation
Let $r_{c.9}$, and $r_{c.3}$ represent the received signals corresponding to the load conditions of the acoustic reflector being at C0.9, and C0.3, respectively. As outlined in Section V-A, the reflection coefficients for the reflector with load impedances $r_{op}$, $r_{sh}$, $r_{c.9}$, and $r_{c.3}$ are $1$, $-1$, $0.9e^{-j\frac{2}{\pi}}$, and $0.3e^{-j\frac{2}{\pi}}$, respectively. Consequently, based on equation (6), we have that
@@
Now, we can use the above ratio to check the accuracy of the quadrature component generated by the reflector. As shown in Fig. 11 (b), the average strengths of differential signals $r_{op}-r_{sh}$, $r_{op}-r_{c.9}$, and $r_{op}-r_{c.3}$ are 31 mV, 23 mV, and 16 mV, respectively; the ratio of these values is 1: 0.74: 0.52, close to the idea case calculated in (8). This validates the ability of the constructed reflector to generate the negative quadrature component for IQ modulation.
@@
As illustrated in Fig. 12, the average strengths of differential signals $r_{l.9}-r_{c.9}$, $r_{l.9}-r_{c.3}$, and $r_{l.3}-r_{c.3}$ are 25 mV, 17 mV, and 6 mV, respectively; the ratio of these values is 1: 0.68: 0.24, close to the idea case calculated in (9). This result validates the ability of the constructed reflector to generate the positive quadrature component for IQ modulation.
VI-C Lake Tests
VI-C1 Experimental setup
We conducted lake experiments to validate the functionality of our UA-RIS system. The control board and the reflecting array are depicted in Fig. 13 (a) and Fig. 13 (b), respectively, while the experimental setup is illustrated in Fig. 13 (c).
In the test, the arbitrary waveform generator emitted a 28 kHz CW signal at 15 dBm. This signal was then amplified using a Fosi M03 audio power amplifier to drive the BTech omnidirectional acoustic transducer. At the receiving end, another BTech transducer positioned 21 m away from the transmitter captured the acoustic signal. The received waveform was initially amplified by a Sound Devices MM-1 microphone amplifier, which provided a 54 dB gain, and then recorded with a PicoScope 4224 portable oscilloscope.
@@ VI-C2 UA-RIS close to sender
The diagrams of the three load configurations, i.e., the coding schemes used in the first test, are depicted in Fig. 14 (a). In these diagrams, the initial loads of the gray and green reflectors are set to open and short-circuits, respectively. The coordinate system and the azimuth beam patterns generated by these three coding schemes are illustrated in Fig. 14 (b).
In Figure 15, the UA-RIS toggles the load of each reflector between short-circuit and open-circuit states every 20 ms, following the initial load coding scheme illustrated in Fig. 14 (a). This configuration results in varying levels of received signal strength, influenced by the beam pattern generated by UA-RIS. Notably, when the azimuth angle of the main beam is set to 90, the reflected waves align directly with the receiver’s direction. This alignment leads to amplitude variations averaging 280 mV and peaking at up to 400 mV as the reflector’s load cycles between the two states. In contrast, when the beam’s azimuth angles are adjusted to 60 and 45, the differential voltage at the receiver decreases markedly to 100 mV and 33 mV, respectively. This reduction is due to the main lobe of the reflected waves deviating from the receiver’s direct path, which diminishes wave superposition at the receiver. Consequently, this lessens the impact of the reflected wave on the received signal strength.
@@
VI-C3 UA-RIS close to receiver
In the second lake test, the transmitter and receiver remained in their original positions, while the UA-RIS was moved to site B, 15 wavelengths behind the receiver. To evaluate the UA-RIS’s performance with IQ modulation, three different load coding schemes were applied. As illustrated in Fig. 16 (a), the initial load settings for the gray, green, purple, and orange reflectors were configured to open, short, L1.0, and C1.0, respectively. The azimuth beam patterns resulting from these coding schemes are shown in Fig. 16 (b), where the main lobes of the reflected beams are directed at 90, 73, and 180, respectively.
@@ VI-C3 UA-RIS close to receiver
In Fig. 17, the UA-RIS alternates the load of each reflector between short-circuit and open-circuit or L10 and C10 every 66 ms, following the initial coding scheme depicted in Fig. 16 (a). From the figure, significant variations in the intensity of the received signal can be observed, which are dependent on the beam pattern generated by the UA-RIS.
@@
VII Conclusion and Future Work

### v2

@@ IV-A2 Circuit Design
The load network comprises three subnetworks: a resistor network, a capacitive network, and an inductive network. The resistor network has only one stage designed to generate the in-phase or antiphase component required for synthetic reflection. Conversely, both the capacitive and inductive networks feature three stages each, dedicated to producing the quadrature component for synthetic reflection.
Resistor network contains an AD8403AR50 programmable potentiometer, whose resistance is managed by the MCU via the serial peripheral interface (SPI) protocol. Let $R_{0}$ represent the resistance of this potentiometer. According to the reflection coefficient presented in (1), when the load connected to the matching circuit is non-reactive and $Z_{0}$ is a real number, $\Gamma$ has no imaginary component. In this circumstance, the magnitude of the reflection coefficient varies between 0 and 1, while the phase shift of the reflected wave remains fixed at 0° for $R_{0}\geq Z_{0}$, or 180° for $R_{0}<Z_{0}$. For instance, if the characteristic impedance of the transmission line is $Z_{L}\!=\!50\,\Omega$, then according to (1), the reflection coefficient $\Gamma$ is 0.33 when $R_{0}\!=\!100\,\Omega$ and --0.33 when $R_{0}\!=\!25\,\Omega$. In both cases, the reflected signals have identical amplitudes but are 180° out of phase.
According to (1), when $Z_{0}$ is a real number, the load impedance must be a complex number to generate the quadrature component necessary for synthetic reflection. In this scenario, it is impossible to continuously adjust the magnitude of the reflection coefficient by simply altering the resistance of the load impedance. To tackle this issue, in our design, the capacitive network includes three stages, designated as $Z_{1}$ to $Z_{3}$ in Fig. 5. Connecting different stages of this network to the matching circuit yields a reflection coefficient with a consistent phase of --90° and three selectable amplitudes increasing stepwise from 0.3 to 0.9. Conversely, $Z_{4}$ to $Z_{6}$ form a three-stage inductive network that produces a reflection coefficient with a consistent phase of 90°, and three optional amplitudes ranging from 0.3 to 0.9.
For example, consider an acoustic signal with a frequency of 28 kHz and a characteristic impedance $Z_{0}\!=\!50\,\Omega$. Suppose a capacitive subnetwork consists of a resistance of 23.5 $\Omega$ and a capacitance of 128.6 nF. The resulting load impedance is calculated as
$\begin{array}[]{lll}Z_{L}\!\!\!\!&=&\!\!\!23.53+\displaystyle\frac{1}{2\pi\!\times\!28\,\text{kHz}\!\times\!128.6\,\text{nF}}\\ \!\!\!\!&=&\!\!\!23.53-44.12j\Omega.\end{array}$ (6)
Substituting (6) into (1), the reflection coefficient is given by
$\Gamma=\displaystyle\frac{Z_{L}-Z_{0}}{Z_{L}+Z_{0}}=\displaystyle\frac{-26.47-44.12j}{73.53-44.12j}\approx-0.6j.$ (7)
In this case, the reflected signal experiences a phase shift of --90° and an amplitude that is 60% of the incident signal.
With the architecture depicted in Fig. 5, the MCU can effectively manipulate both the phase and magnitude of the reflection coefficient by adjusting the resistance of the potentiometer and the ON/OFF status of the NMOS transistors in each stage of the load network. In the following section, we will introduce the selection of $Z_{0}$ and discuss the strategy for managing the high dynamic range of the acoustic signal in the circuit design.
IV-B High-Impedance Matching
@@
In order to precisely control the phase and amplitude of the quadrature component for accurate synthetic reflection, the wiper resistance of programmable potentiometers cannot be ignored. For instance, models such as the AD8403 [29] and AD5245 [30] typically exhibit a wiper resistance of about 50 $\Omega$, while the MAX5427 has a resistance of 100 $\Omega$ [31]. Consequently, the minimum resistance of the load network in (1) does not approach zero but remains significantly higher. In scenarios where the impedance of the reflector is matched to 50 $\Omega$ (i.e., $Z_{0}\!=\!50$$\Omega$), the reflection coefficient consistently remains positive. As a result, the reflector will be incapable of generating an antiphase wave.
@@
IV-C Highly Dynamic Incident Waves
@@
V Performance Evaluation
This section evaluates the performance of the proposed UA-RIS through experiments. Initially, we examine the power consumption of the system. Subsequently, the results collected from COMSOL simulation, as well as controlled tank tests and open-water lake experiments are presented to demonstrate the viability of the UA-RIS in real-world applications.
V-A Power Consumption
Although the UA-RIS does not transmit acoustic signals, it still requires an energy supply to maintain the operation of active components for synthetic reflection. In our implementation, the power consumption of the system primarily originates from four components: the MCU, twelve I/O extenders, and six programmable potentiometers. In this section, we thoroughly analyze the power consumption of UA-RIS operating in various states.
@@ V-A Power Consumption
V-B COMSOL Simulation Results
To assess the performance of the UA-RIS, simulations are carried out using COMSOL Multiphysics. The 3D structure of the acoustic reflector is illustrated in Fig. 7. The head mass is constructed from aluminum, while the tail mass and prestressing rod are made of AISI 4340 steel. PZT-4 is employed as the acoustic vibrational element.
Figure 7: 3D structure of the acoustic reflector constructed for the simulation.
In the simulation, a monochromatic plane wave at 28 kHz is used as the signal source. To enhance simulation accuracy, a five-step procedure is followed. First, the impedance of the PZT disk is determined by measuring the ratio of open-circuit voltage to short-circuit current. Based on this impedance value, appropriate matching circuit is constructed. A load is then introduced, and the resulting voltage response under the plane wave excitation is recorded. This measured voltage, together with the reflection coefficients of the load, is used to compute the reflected voltage. Finally, with the incident wave deactivated, the computed reflected voltage is applied to the matching circuits, and the resulting pressure waveform is recorded to represent the signal reflected by the structure under the specified load conditions.
Figure 8: Stress distributions of acoustic reflectors under various load conditions measured at 112 $\mu$s.
Fig. 8 illustrates the stress distribution of acoustic reflectors measured under four distinct load conditions. It is evident that the stress profiles vary significantly across these conditions. As shown in the figure, at 112 $\mu$s, the stress in the reflector with an open-circuit load reaches its negative peak, whereas the stress in the reflector with a short-circuit load reaches its positive peak. This observation is consistent with the theoretical analysis presented in Section III-A, which indicates that reflector coefficients generated by open and short circuits have identical amplitudes but opposite phases. In contrast, the stress in reflectors loaded with L0.9 and C0.9 is nearly zero at this time instant. This is because the reflector coefficients associated with inductive and capacitive loads exhibit phase shifts of 90 and --90, respectively, relative to the open circuit. Consequently, when the stress for the open-circuit load reaches its peak at 112 $\mu$s, the stress for the inductive and capacitive load conditions crosses zero.
Figure 9: Reflected beams generated by a UA-RIS with 8 acoustic reflectors using the synthetic reflection scheme.
Fig. 9 illustrates the reflected beams generated by a UA-RIS comprising eight reflectors under a monochromatic plane wave excitation. The simulation environment is a water tank with dimensions of 1 m×$\times$1 m×$\times$2 m, and the source signal operates at a frequency of 28 kHz. Each reflector is labeled with a red identification number, and the spacing between adjacent reflectors is set to $2\lambda$, where $\lambda$ denotes the acoustic wavelength. To evaluate the acoustic directivity, 72 probes are uniformly placed along the blue dotted line on the Y-Z plane, each positioned 75 cm from the center of the reflector array. Pairs of reflectors aligned along the X-axis are grouped for synthetic reflection, and their respective load impedances are provided in Table II. With this configuration, the UA-RIS is designed to produce a directional beam with maximum gain at 225 in the Y-Z plane.
Table II: Load impedance for different methods.
ID 1 2 3 4 5 6 7 8
Syn. 150 $\Omega$ 150 $\Omega$ 667 $\Omega$ C06 18 $\Omega$ L10 50 $\Omega$ C10
1-bit Open Open Open Open Short Short Open Open
As shown in Fig. 9, the reflected beam exhibits strong directivity at 225 as intended. Additionally, a grating lobe is observed near 240, resulting from the reflector spacing exceeding $1\lambda$. In the simulation, the absence of matching materials between the head mass and the surrounding water leads to a significant acoustic impedance mismatch at the interface, thereby reducing the amplitude of the reflected pressure signal. In practical implementations, this limitation can be addressed by applying multiple matching layers, an approach commonly employed in commercial acoustic transducers, to effectively minimize reflection losses at exposed boundaries.
Figure 10: Normalized reflected beam patterns produced by different schemes.
In Fig. 10, we compare the reflected beam patterns generated by the proposed synthetic scheme and the conventional 1-bit coding scheme that is commonly used in RF-RIS. The beam amplitudes are normalized to their respective peak values. The simulation setup is identical to that in Fig. 9, and the load impedances applied to each reflector are listed in Table II.
As shown in Fig. 10, both schemes successfully steer the main beam toward 225 as expected. However, the synthetic scheme produces a noticeably narrower main lobe compared to the 1-bit coding approach, indicating improved beam directivity. Moreover, the side lobe levels produced by the synthetic scheme are significantly lower. For example, at 265, 275, and 290, the normalized amplitudes of side lobes under the synthetic scheme are 0.46, 0.43, and 0.27, respectively, whereas the corresponding values under the 1-bit scheme are 0.79, 0.72, and 0.48. These results demonstrate that the UA-RIS configured with the synthetic scheme achieves superior beamforming performance with enhanced directivity and reduced side lobe levels.
V-C Tank Experiments
V-C1 Experimental settings
@@
V-C2 Resistive load modulation
@@
V-C3 Reactive load modulation
The multipath model presented in (8) can also be used to validate the reflector’s ability to produce the quadrature component for synthetic reflection. This can be achieved by programming the MCU to sequentially connect the subnetwork shown in Fig. 5 with the acoustic reflector. The waveform of the received signal is depicted in Fig. 15 (a) [^1] The experiment in Fig. 15 was conducted on a different day than those in Figs. 12 to 14. This difference resulted in varying acoustic environments and channel states, indicating that this experiment should be considered independent., where the source is 28.23 kHz CW with 20 V peak-to-peak voltage. In the figure, labels such as C0.3 and L0.9 indicate that the acoustic reflector is connected to the capacitive and inductive load network with reflection coefficients of 0.3 and 0.9 amplitude, respectively.
@@ V-C3 Reactive load modulation
Let $r_{c.9}$, and $r_{c.3}$ represent the received signals corresponding to the load conditions of the acoustic reflector being at C0.9, and C0.3, respectively. As outlined in Section IV-A, the reflection coefficients for the reflector with load impedances $r_{op}$, $r_{sh}$, $r_{c.9}$, and $r_{c.3}$ are $1$, $-1$, $0.9e^{-j\frac{2}{\pi}}$, and $0.3e^{-j\frac{2}{\pi}}$, respectively. Consequently, based on (8), we have that
@@
Now, we can use the above ratio to check the accuracy of the quadrature component generated by the reflector. As shown in Fig. 15 (b), the average strengths of differential signals $r_{op}-r_{sh}$, $r_{op}-r_{c.9}$, and $r_{op}-r_{c.3}$ are 31 mV, 23 mV, and 16 mV, respectively; the ratio of these values is 1: 0.74: 0.52, close to the idea case calculated in (10). This validates the ability of the constructed reflector to generate the negative quadrature component for synthetic reflection.
@@
As illustrated in Fig. 16, the average strengths of differential signals $r_{l.9}-r_{c.9}$, $r_{l.9}-r_{c.3}$, and $r_{l.3}-r_{c.3}$ are 25 mV, 17 mV, and 6 mV, respectively; the ratio of these values is 1: 0.68: 0.24, close to the idea case calculated in (11). This result validates the ability of the constructed reflector to generate the positive quadrature component for synthetic reflection.
In the tank tests, the relatively strong reflections caused by the tightly spaced setup may lead to an overestimation of reflector gains in more complex or challenging environments. To address this limitation, lake tests are conducted in Section V-D to further evaluate the performance of UA-RIS in a real-world scenario with dynamic channel conditions.
V-D Lake Tests
V-D1 Experimental setup
We conducted lake experiments to validate the functionality of our UA-RIS system. The control board and the reflecting array are depicted in Fig. 17 (a) and Fig. 17 (b), respectively, while the experimental setup is illustrated in Fig. 17 (c). In the test, the arbitrary waveform generator emitted a 28 kHz CW signal at 15 dBm. This signal was then amplified using a Fosi M03 audio power amplifier to drive the BTech omnidirectional acoustic transducer. At the receiving end, another BTech transducer positioned 21 m away from the transmitter captured the acoustic signal. The received waveform was initially amplified by a Sound Devices MM-1 microphone amplifier, which provided a 54 dB gain, and then recorded with a PicoScope 4224 portable oscilloscope.
@@ V-D2 UA-RIS close to sender
The diagrams of the three load configurations, i.e., the coding schemes used in the first test, are depicted in Fig. 18 (a). In these diagrams, the initial loads of the gray and green reflectors are set to open and short-circuits, respectively. The coordinate system and the azimuth beam patterns generated by these three coding schemes are illustrated in Fig. 18 (b). In the figure, the beam patterns are numerically derived by processing the reflection patterns corresponding to different RIS configurations. Each configuration corresponds to a fixed incident angle ($90^{\circ}$, $60^{\circ}$, and $45^{\circ}$), and the reflected field is synthesized by correlating the phase-coded reflection sequences with theoretical steering vectors over a range of azimuth angles in the X-Y plane. The resulting beam patterns shown in Fig. 18 (b) are computed based on this correlation to indicate the directional gain for each configuration.
In Fig. 19, the UA-RIS toggles the load of each reflector between short-circuit and open-circuit states every 20 ms, following the initial load coding scheme illustrated in Fig. 18 (a). This configuration results in varying levels of received signal strength, influenced by the beam pattern generated by UA-RIS. Notably, when the azimuth angle of the main beam is set to 90, the reflected waves align directly with the receiver’s direction. This alignment leads to amplitude variations averaging 280 mV and peaking at up to 400 mV as the reflector’s load cycles between the two states. In contrast, when the beam’s azimuth angles are adjusted to 60 and 45, the differential voltage at the receiver decreases markedly to 100 mV and 33 mV, respectively. This reduction is due to the main lobe of the reflected waves deviating from the receiver’s direct path, which diminishes wave superposition at the receiver. Consequently, this lessens the impact of the reflected wave on the received signal strength.
@@
V-D3 UA-RIS close to receiver
In the second lake test, the transmitter and receiver remained in their original positions, while the UA-RIS was moved to site B, 15 wavelengths behind the receiver. To evaluate the UA-RIS’s performance with synthetic reflection, three different load coding schemes were applied. As illustrated in Fig. 20 (a), the initial load settings for the gray, green, purple, and orange reflectors were configured to open, short, L0.9, and C0.9, respectively. The azimuth beam patterns resulting from these coding schemes are shown in Fig. 20 (b), where the main lobes of the reflected beams are directed at 90, 73, and 180, respectively.
@@ V-D3 UA-RIS close to receiver
In Fig. 21, the UA-RIS alternates the load of each reflector between short-circuit and open-circuit or L0.9 and C0.9 every 66 ms, following the initial coding scheme depicted in Fig. 20 (a). From the figure, significant variations in the intensity of the received signal can be observed, which are dependent on the beam pattern generated by the UA-RIS.
@@
V-D4 Discussions
In the lake tests, using CW signals primarily validate steady-state characteristics such as beam steering and reflection capabilities of UA-RIS, which served as foundational assessments for this work. Practical modulation schemes that commonly used for underwater acoustic communications, such as frequency shift keying (FSK), phase shift keying, and orthogonal frequency division multiplexing (OFDM) will introduce additional complexities and transient behaviors not fully captured by CW testing. However, preliminary analytical and experimental results collected in the field test provide strong evidence that performance improvements are expected when transitioning to practical modulation schemes.
For precise beamforming, the UA-RIS needs to be aware of its own attitude, which may vary due to ocean currents. To ensure optimal performance, real-time attitude estimation can be achieved by incorporating data from an inertial measurement unit (IMU) and a pressure sensor. For instance, the gyroscope within the IMU provides angular velocity measurements that enable tracking of three-dimensional orientation changes over time, while the pressure sensor supplies depth information essential for vertical positioning.
Our proof-of-concept experiments are based on a fixed, known geometry and therefore do not incorporate incident-angle estimation or transceiver localization. Accurate estimation of the angle of incident waves for steering the reflected signal in the desired direction is essential yet challenging tasks. To solve this problem, UA-RIS can leverage the spatial characteristics of the received signal across the reflection elements to estimate the angle of arrival (AoA) of an incident signal. In scenarios where the RIS is equipped with sensing capabilities either through embedded low-power sensors or via hybrid active elements, it becomes possible to treat the RIS as an antenna array that samples the incoming wavefront. By modeling the received signal as a superposition of array response vectors parameterized by the incident angle, traditional array signal processing techniques such as MUSIC (multiple signal classification) or ESPRIT (estimation of signal parameters via rotational invariance techniques) can be applied to estimate the AoA with high resolution. This approach allows precise directional information to be extracted, which is critical for dynamically configuring the RIS reflection pattern to optimize signal propagation and enhance communication performance.
VI Conclusion and Future Work